### v2

# Exact Simulation of Longitudinal Data from Marginal Structural Models

Abstract
Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and clinical trial design. While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, it is difficult to ensure alignment between conditional distributions and the target marginal causal effects. This misalignment presents a fundamental challenge in simulating data that adheres to marginal structural model specifications. To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model. Recognizing the importance of time-to-event outcomes in clinical research, we extend our approach to accommodate survival models. Compared to existing approaches, our method offers several advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids restrictive assumptions about the data-generating process; and remains computationally efficient by requiring only the evaluation of analytic functions. This last benefit contrasts with methods that use computationally intensive techniques such as Monte Carlo approximations or numerical integration. Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility. Our method will facilitate researchers in effectively simulating data with target causal structures for their specific scenarios.
1 Introduction
Longitudinal data are fundamental for studies analyzing the causal effect of a treatment or exposure on an outcome, especially for research into chronic diseases such as diabetes or HIV. Unlike cross-sectional studies, longitudinal designs capture the dynamic nature of treatment regimens, which often vary over time due to clinical decisions, patient adherence, or external factors. Even in clinical trials with strict adherence protocols, intercurrent events—such as treatment discontinuation or supplementary medication use—are common. By accounting for these treatment dynamics, longitudinal data enable the estimation of meaningful causal quantities, such as the “on-treatment” estimand, which quantifies the treatment effect assuming treatment is administered as intended.
A central challenge in analyzing longitudinal data is addressing time-varying confounding, where a time-dependent covariate predicts both the outcome and future treatments, while itself being influenced by past treatments. For example, in the case study in Robins et al. (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression. The presence of such time-varying confounding renders conventional adjustment methods, like regression and stratification, generally biased.
Marginal structural models, introduced by Robins (1997), have emerged as a powerful tool for estimating causal effects under time-varying confounding. Model parameters can be consistently estimated using inverse probability of treatment weighting (Robins, 1999, 2000). Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), marginal structural models have gained prominence due to their interpretability and ease of implementation. A systematic review by Clare et al. (2019) highlights their widespread use in applied research.
As causal inference methods for longitudinal data are continually proposed to address complex real-world settings, simulation studies are essential to evaluate their effectiveness for practical use. These studies assess important properties such as bias, confidence interval coverage, relative efficiency, and robustness to violations of the assumptions used, guiding both methodological development and application. Sample size calculations also benefit from simulations. Prior to conducting a trial or an observational study, it is crucial to determine the number of participants or number of observational units required to achieve adequate statistical power. While standard formulas can be used for time-fixed treatments, more complex scenarios with time-varying treatments make deriving analytical formulas impractical or impossible. In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
A fundamental requirement for both method evaluation and sample size determination is the ability to simulate data that conform to a specified causal model. However, simulating longitudinal data from a pre-specified marginal structural model is not a straightforward task. While generating data sequentially from structural equations—creating covariates, treatment status, and then the outcome conditioned on past information—may seem intuitive, these models rely on conditional rather than marginal specifications. This makes it difficult to sequentially simulate with conditional distributions in a way that aligns with a pre-specified marginal model; in particular, it is extremely challenging to sample from a null hypothesis —the so-called g-null paradox (Robins and Wasserman, 1997).
The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models. Section 3 introduces key preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas. In Section 4, we introduce our proposed methodology and illustrate it in Section 4.4 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004). Section 5 extends the simulation approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, including an illustrative example. The paper concludes with a discussion in Section 6. Example code for the simulations in Sections 4.4 and 5.3 is available at https://github.com/XiLinStats/LongSiMSM.
2 Related Literature
As noted in Section 1, the challenge lies in ensuring that the simulated data adhere to a specified marginal model, rather than simply following sequential conditional distributions in a structural equation model. Existing approaches address this issue through two strategies:
  1. (i)
    Simulating from conditional models and deriving or approximating the corresponding marginal models post-hoc.
  2. (ii)
    Bypassing conditional simulations and simulating directly from the target marginal model.
Several studies have explored the first approach. Xiao et al. (2010) used sequential conditional models to approximate Cox marginal structural models under the rare disease assumption, i.e. very low failure rates, which mitigates the impact of noncollapsibility but assumes the absence of unmeasured confounding between time-varying confounders and outcomes. Young and Tchetgen Tchetgen (2014) employed the g-formula to bridge conditional and marginal models in discrete-time Cox marginal structural models. While theoretically feasible for discrete time-varying covariates, deriving marginal models becomes intractable for continuous covariates due to the need for numerical integration. Simplifications, such as the Markov property or rare disease assumptions, are often necessary but limit generalizability. Sofrygin et al. (2017) developed the simcausal R package for flexible simulation from non-parametric structural equation models. The workflow involves specifying structural equations and sequentially simulating data from conditional distributions. However, deriving analytic expressions for marginal causal quantities under this conditional parameterization is often intractable and prone to model misspecification. To address this, users typically specify low-dimensional working marginal structural models (Neugebauer and van der Laan, 2007) and estimate their parameters via Monte Carlo simulations of counterfactual data. This introduces two limitations: first, the true marginal model remains unknown; second, reverse-engineering simulations to match a target causal function is challenging. Keogh et al. (2021) described simulation from additive-hazard marginal structural models (Aalen, 1989). They utilized the result that if the conditional hazard model follows an additive form, the equivalent marginal hazard model will also be additive. While the marginal model retains the correct form, true parameter values are generally intractable and require Monte Carlo approximations. A further limitation is the difficulty in ensuring non-negative hazards for continuous covariates.
Alternative methods focus on simulating directly from the target marginal model. Havercroft and Didelez (2012) simulated directly from a discrete Cox marginal structural model, introducing time-varying confounding indirectly via a shared latent variable. However, this approach assumes no direct effect of time-varying confounders on outcomes, which may not hold in practice, e.g. the effect of CD4 count on HIV survival (Robins et al., 2000). Evans and Didelez (2024) proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly, enabling direct simulation from the exact marginal model while explicitly modeling time-varying covariates. This addresses key limitations of earlier methods. Building on this, Seaman and Keogh (2024) applied the frugal parameterization to simulate survival outcomes, using a one-dimensional “risk score” to summarize multi-dimensional covariates. While simplifying dependence structures, this method relies on computationally expensive Monte Carlo approximations to estimate the cumulative distribution functions of the risk score, making it infeasible for continuous treatments or baseline variables.
The aims for a new simulation approach therefore include the ability to: precisely specify the marginal structural model; include multiple time-varying covariates, treatments, outcomes, and censoring; have good control over their distributions; avoid restrictive assumptions about the data-generating process. In addition, we would like it to scale well in the number of time-points and variables, and be applicable to real-world problems.
3 Preliminaries
3.1 The frugal parameterization
As discussed in the previous section, the fundamental challenge in simulating data from longitudinal marginal structural models arises from the fact that, while the joint data distribution can be easily factorized into conditionals for simulation, the causal quantity of interest is a marginal distribution. Earlier methods, as detailed in Section 2, tackled this by deriving the marginal distribution implied by the conditional factorization. However, Evans and Didelez (2024) advanced this area with the introduction of the frugal parameterization, which explicitly parameterizes the causal marginal and constructs the rest of the joint distribution around it. This development has made the simulation method presented in this paper possible.
The frugal parameterization decomposes the joint distribution $p(\bm{l},a,y)$ into three separate pieces:
  1. (i)
    the past: $p_{\bm{L}A}(\bm{l},a)$—the joint distribution of covariates and treatment prior to observing outcomes;
  2. (ii)
    the causal margin: $p^{*}_{Y\mid A}(y\,|\,do(a))$—the causal margin of interest, set according to the marginal structural models relevant to the causal question;
  3. (iii)
    a dependency measure: $\phi^{*}_{Y\bm{L}\mid A}$—captures the dependency between the outcome $Y$ and the covariates $\bm{L}$ under intervention.
A common choice of this dependency measure is a copula, which we elaborate on in the following subsection.
3.2 Copulas
We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure. This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$. According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as
$F(x_{1},\ldots,x_{d})=C\left\{F_{1}(x_{1}),\ldots,F_{d}(x_{d})\right\}=C(u_{1},\ldots,u_{d}),$
where $\bm{X}=(X_{1},\ldots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\ldots,d$. The copula then takes the form
$C(u_{1},\ldots,u_{d})=F\{F_{1}^{-1}(u_{1}),\ldots,F_{d}^{-1}(u_{d})\},\qquad(u_{1},\ldots,u_{d})\in[0,1]^{d}.$
The choice of copula can be considered independently of the univariate margins.
3.2.1 Pair-copula construction
While there exist a wide variety of parametric bivariate copulas, the options for higher-dimensional copulas are limited, with Gaussian and Student t-copulas being common choices. Extending symmetric bivariate Archimedean copulas to multivariate settings is also possible but imposes strong restrictions, such as requiring the same level of dependence among all possible pairs. Building on the work of Joe (1996), Bedford and Cooke (2001, 2002) introduced pair-copula constructions, a flexible and intuitive approach to constructing multivariate copulas by modeling multivariate data as a series of bivariate copulas. This hierarchical structure incorporates additional variables at each level, using pair-copulas as the fundamental building blocks.
Following the general formula given in Joe (1997), the conditional distribution of one variable $y$ given $d$ other variables can be decomposed to $d$ bivariate copula density functions, each of which can be specified flexibly:
$f(y\,|\,\bm{x})=c_{YX_{j}|\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}\cdot f(y\,|\,\bm{x}_{-j}),$ (1)
where $\bm{x}$ is a $d$-dimensional vector; $x_{j}$ is an arbitrarily chosen component of $\bm{x}$, and $x_{-j}:=(x_{1},\ldots,x_{j-1},x_{j+1},\ldots,x_{d})$ denotes $\bm{x}$ with its $j$th entry removed. Applying (1) recursively to the conditional densities, we have:
$f(y\,|\,\bm{x})$ $=f(y)\cdot\prod_{i=1}^{d}c_{YX_{d+1-i}|\bm{X}_{1:{d-i}}}\{F(y\,|\,\bm{x}_{1:{d-i}}),F(x_{d+1-i}\,|\,\bm{x}_{1:{d-i}})\},$ (2)
where each copula in the product in (2) is bivariate; this defines a pair-copula. In principle, any ordering of the conditioned variables $x_{1},x_{2},\ldots,x_{d}$ can be selected, although some orderings may be easier to use than others. In our proposed method, we choose an ordering that is convenient and efficient for simulation.
In conclusion, under appropriate regularity conditions, a multivariate density can be expressed as a product of pair-copulas, acting on different conditional distributions.
3.2.2 Simulating from pair-copula constructions
The pair-copula construction includes conditional distributions of the form $F(y\,|\,\bm{x})$. Joe (1996) demonstrated the following result:
$F(y\,|\,\bm{x})=\frac{\partial C_{YX_{j}\mid\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}}{\partial F(x_{j}\,|\,\bm{x}_{-j})}.$ (3)
For notational convenience, we introduce the h-function $h(u_{1},u_{2};\Theta)$ from Aas et al. (2009). Let $u_{1}$ and $u_{2}$ be random variables with a uniform distribution on $(0,1)$. The $h$-function is defined as
$h(u_{1},u_{2};\Theta):=F(u_{1}\,|\,u_{2})=\frac{\partial}{\partial u_{2}}C_{12}(u_{1},u_{2};\Theta)$
where the second argument in $h(\cdot;\Theta)$ is the conditional variable, and $\Theta$ represents the set of copula parameters. The $h$-function is effectively the conditional copula $C(u_{1}\,|\,u_{2};\Theta)$.
Remark 3.1.
The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward by definition for parametric families. Examples include Gaussian, Student t, Clayton, Frank, and Gumbel copulas.
The $h$-functions are foundational for simulating from $F(y\,|\,\bm{x})$. Similar to (2), by imposing an ordering $1,\ldots,d$ on the components of $\bm{x}$, we can recursively apply $h$-functions to obtain:
$F(y\,|\,\bm{x})$ $=h\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}}),\Theta_{d}\}$
$=h[h\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}};\Theta_{d-1}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{d}]$
$\ldots$
$=h[\ldots h\{h(F(y),F(x_{1});\Theta_{1}),F(x_{2}\,|\,x_{1});\Theta_{2}\},\ldots,F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{d}].$ (4)
Intuitively, the bivariate copulas “weave” the dependence between the outcome and each of the covariates. Thus, given a value for $\bm{X}$ we can simulate $Y$ by first drawing a uniform $U\sim{\rm U}(0,1)$, and then evaluating:
$F(Y)=h^{-1}[\ldots h^{-1}\{h^{-1}(U,\,F(X_{1});\Theta_{1}),F(X_{2}\,|\,X_{1});\Theta_{2}\},\ldots,F(X_{d}\,|\,X_{1},\ldots,X_{d-1});\Theta_{d}].$ (5)
This means that to simulate $Y$ given a $d$-dimensional vector $\bm{X}$, we recursively calculate the $d$ inverse $h$-functions, forming a pair-copula construction. This underpins our proposed simulation method.
4 Simulation Algorithm
[…]
Figure 1: An example directed acyclic graph showing temporal relationships in a longitudinal discrete setting, with baseline covariates $Z$; treatments $A_{0},\ldots,A_{K}$; time-varying covariates $L_{0},\ldots,L_{K}$ and an outcome $Y$ observed at the end of the K follow up intervals. Given the topological ordering $Z,L_{0},A_{0},L_{1},A_{1}\ldots,L_{K},A_{K},Y$, any variable may depend upon any subset of the previous variables.
We first consider a longitudinal study with $n$ individuals observed at discrete time points, $\mathcal{K}=\{0,\ldots,K\}$. At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded. Additionally, $Z$ represents a set of static baseline covariatesatent variables, if present, are written as $\{H_{k}\}_{k\in\mathcal{K}}$. To indicate the history of a variable up to time $k$, we use bar notation. For instance, $\bar{A}_{k}=\{A_{0},\ldots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},\ldots,L_{k}\}$, with the convention $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$. The outcome $Y$ is observed at the final time point $K$. Figure 1 is a graphical representation of this setup, which is common in medical and epidemiological research. Examples include examining the effects of an HIV treatment on infection status assessed at the end of the study period (Robins et al., 2000), and assessing the impact of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004). In Section 4.4, we illustrate our method through a simulation study that replicates such settings.
Following the notation of Evans and Didelez (2024), we use an asterisk to represent causal or interventional distributions where treatment is set to a specific value $A_{k}=a$ via intervention. In graphical terms, this corresponds to removing all arrows directed into $A_{k}$.
4.1 The fully-conditional approach
To outline our approach, we initially assume the time-varying covariate $L_{k}$ is one-dimensional and omit any latent process $\{H_{k}\}_{k\in\mathcal{K}}$; the extension to multi-dimensional time-varying covariates and latent variables is discussed in Appendix A.
The standard method for simulation of longitudinal data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ is by time step in blocks using the Rosenblatt transformation (Rosenblatt, 1952). Specifically:
  1. 1.
    Baseline simulation: Simulate static baseline variables, $z$ following a specified factorization, e.g.  $F_{Z_{1}},F_{Z_{2}\mid Z_{1}},\ldots$.
  2. 2.
    Time-step simulation: At each time point $k=0,\dots,K$, simulate $l_{k}=F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}^{-1}(w_{k,1},\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$ and $a_{k}=F_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{l}_{k},\bar{a}_{k-1})$, where $w_{k,1},w_{k,2}\sim\operatorname{U}(0,1)$.
  3. 3.
    Outcome simulation: Given the simulated covariates and treatments, generate the outcome $y=F_{Y\mid Z,\bar{L}_{K},\bar{A}_{K}}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K})$, with $w_{Y}\sim\operatorname{U}(0,1)$.
Our proposed method diverges from this fully-conditional approach at Step 3. Instead of specifying the conditional distribution $F_{Y\mid Z\bar{L}_{K}\bar{A}_{K}}$—which inherently introduces misalignment between conditional and the desired marginal models—we introduce a margin-preserving parameterization for outcome simulation.
4.2 Frugal parameterization for outcome simulation
Our method uses the frugal parameterization to simulate outcomes, avoiding potentially intractable analytical calculations to obtain the marginal structural model. Evans and Didelez (2024) showed that the fully-conditional distribution of the outcome can be frugally parameterized as:
$p_{Y\mid\bar{L}_{K},\bar{A}_{K}}(y\,|\,\bar{l}_{k},\bar{a}_{K})=p^{*}_{Y\mid\bar{A}_{K}}(y\,|\,do(\bar{a}_{K}))\cdot c^{*}_{Y\bar{L}_{K}\mid\bar{A}_{K}}(u^{*}_{Y},u_{L_{0}},\dots,u_{L_{K}}\,|\,do(\bar{a}_{K})),$ (6)
where $u^{*}_{Y}=F_{Y\mid\bar{A}_{K}}(y\,|\,do(\bar{a}_{K}))$ and $u_{L_{k}}=F_{L_{k}\mid\bar{A}_{K}}(l_{k}\,|\,do(\bar{a}_{K}))$ for $k=0,\ldots,K$ are the marginal quantiles under interventions that set $\bar{A}_{K}=\bar{a}_{K}$.
Remark 4.1.
For brevity, we omit conditioning on baseline variables $Z$ in all quantities in (6), treating them as if no baseline variables exist. This omission will continue hereafter, though all expressions can be extended to condition on $Z$ by including it in their respective conditioning sets.
This decomposition implies that we can simulate $Y$ through the multivariate copula and its interventional margin. Applying the pair-copula construction in (2) to (6), we obtain:
$p(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=p(y\,|\,do(\bar{a}_{K}))\cdot\prod_{k=0}^{K}c^{*}\{F(y\,|\,\bar{l}_{k-1},do(\bar{a}_{K})),F(l_{k}\,|\,\bar{l}_{k-1},do(\bar{a}_{K}));\Theta_{k}\}.$ (7)
Note that we omit some subscripts on densities for conciseness where doing so does not introduce ambiguity. This decomposes the conditional density into the product of $K+1$ pair-copulas and the target marginal causal density. The pair-copula linking $Y$ and $L_{k}$ is parameterized by $\Theta_{k}$, which may depend on intervened treatments $do(\bar{A}_{k})$, past time-varying confounders $\bar{L}_{k-1}$ and baseline covariates $\bm{Z}$. Under intervention, graphically speaking, all arrows into nodes $\bar{A}_{k}$ are removed in the directed acyclic graph, rendering $L_{k}$ independent of subsequent interventions. Consequently, $F(l_{k}\,|\,\bar{l}_{k-1},do(\bar{a}_{K}))=F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$ for $k=1,\ldots,K$ and $F(l_{0}\,|\,do(\bar{a}_{K}))=F(l_{0})$. Following the simulation strategy from pair-copula constructions outlined in Section 3.2.2, the fully-conditional distribution of the outcome is alternatively parameterized as:
$F(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=h\big{[}\dots h\{h(F^{*}(y\,|\,do(\bar{a}_{K})),F(l_{0});\Theta_{0}),F(l_{1}\,|\,l_{0},a_{0});\Theta_{1}\}\dots;\Theta_{K}\big{]},$
which is centered around the causal margin $F^{*}(y\,|\,\bar{a}_{K})$.
Through inversion the outcome can be simulated in two steps:
  1. 1.
    Calculate the marginal quantile of the outcome, $u^{*}_{Y}$, as:
    $u^{*}_{Y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{Y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1});\Theta_{K}),$
    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2});\Theta_{K-1}\big{\}}\dots,F(l_{0});\Theta_{0}\big{]}$ (8)
    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\ldots,w_{0,1};\Theta_{0}\big{]},$
    where $w_{Y}\sim\operatorname{U}(0,1)$.
  2. 2.
    Obtain $y$ through the inverse cumulative distribution function of the marginal distribution: $y=F^{*-1}_{Y\mid\bar{A}_{K}}(u^{*}_{Y}\,|\,do(\bar{a}_{K}))$.
An intuitive way to understand the calculation in Step 1 is that the inverse $h$-functions “disentangle” the dependencies between $Y$ and each of $L_{K},L_{K-1},\dots L_{0}$ sequentially, and remove the conditioning of $Y$ on $\bar{L}_{K}$ until the marginal quantile of $Y$ is obtained. The quantiles in the inverse $h$-functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$, correspond exactly to the random standard uniform variables $w_{k,1}$ used to simulate $l_{k}$ at time step $k$. This enables direct substitution of the $l_{k}$ quantiles into (8) without additional computation.
4.3 Simulation Algorithm
In summary, to simulate the data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$, the user specifies:
  • the population distribution of the static baseline covariates $Z$: $F_{Z}$, either as a joint distribution, or through a conditional factorization;
  • the development of the time-varying confounder $L_{k}$ as influenced by $\bm{Z}$ and $\bar{A}_{k-1}$: $F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}$;
  • treatment assignment based on the time-varying confounder $\bar{L}_{k}$, past treatment and baseline covariates $Z$: $F_{A_{k}\mid\bm{Z},\bar{L}_{k},\bar{A}_{k-1}}$;
  • the target causal marginal structural model: $F_{Y\mid\bar{A}_{k},Z}^{*}(\cdot\,|\,do(\bar{a}_{K}),z)$;
  • $K+1$ bivariate copulas describing dependencies between outcome $Y$ and $L_{0},\ldots,L_{K}$:
    $C_{YL_{k}\mid Z,do(\bar{A}_{k}),\bar{L}_{k-1}}(\cdot;\Theta_{k})\quad\text{for}\quad k=0,\ldots,K.$
Section 4.4 provides an example of specifying and interpreting these inputs through a simulation study that captures real-world complexities.
Algorithm 1 below details the simulation procedure:
Algorithm 1 Simulation algorithm for an individual $i$ over $K$ follow-ups
Sample $w_{z}\sim U(0,1)$
$z\leftarrow F_{Z}^{-1}(w_{z})$
Let $k=0$
while $k\leq K$ do
     Sample $w_{k,1},w_{k,2}$ independent standard uniform variables
     $l_{k}\leftarrow F^{-1}_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}(w_{k,1}\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$
     $a_{k}\leftarrow F^{-1}_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}(w_{k,2}\,|\,z,\bar{l}_{k},\bar{a}_{k-1})$
     $k\leftarrow k+1$
end while
Sample $w_{y}\sim U(0,1)$
Let $\nu_{0}=w_{y}$
for $j=0$ to $K$ do
     $\nu_{j+1}\leftarrow h^{-1}(\nu_{j},w_{(K-j),1};\Theta_{K-j})$
end for
$u^{*}_{y}\leftarrow\nu_{K+1}$
$y\leftarrow F_{Y\mid\bar{A}_{k},Z}^{*-1}(u^{*}_{y}\,|\,do(\bar{a}_{K}),z)$
return $(z,\bar{l}_{K},\bar{a}_{K},y)$
Extending the simulation algorithm to include a latent process $\{H_{k}\}_{k\in\mathcal{K}}$, or multiple time-varying confounders is straightforward. The key modification lies in the outcome simulation: the marginal quantile of $Y$ is obtained by sequentially removing the conditioning on all the time-dependent confounders at each step, through inverse $h$-functions. We provide an illustrative example including a latent variable $H_{k}$ in Appendix A.
4.4 Illustrative example
We demonstrate our simulation method by replicating the setting in Bodnar et al. (2004), which examines the causal effect of iron supplement use during pregnancy on the odds of anemia at delivery. Participants are randomized into treatment and control groups and followed up over four visits leading up to delivery. Treatment is not time-fixed as participants are assumed to adhere to assigned treatment up to $k=2$, after which standard clinical protocols dictate further treatment based on hemoglobin levels.
[…]
$\mathbb{P}[Y=1\,|\,do(\bar{A}_{4}),B,C]=\operatorname{expit}(-2+0.1\,B+0.02\,C-\beta\,\operatorname{cum}(\bar{A}_{4})),$
where $\operatorname{cum}(\bar{A}_{4})=\Sigma_{i=1}^{4}A_{i}$. Here, $\exp(\beta)$ represents the causal odds ratio for an incremental dose, and $\exp(4\beta)$ corresponds to the “always treated” efficacy estimand.
We use a t-copula $T_{\rho,\operatorname{df}}$ (with $\rho=-0.5$ and $\operatorname{df}=5$) across all pair-copula construction components to model the dependency between $Y$ and each of $L_{0},L_{1},L_{2},L_{3}$, and $L_{4}$. Compared to a Gaussian copula, the t-copula captures heavier-tailed dependency, capturing stronger associations in the tails reflecting stronger associations in extreme scenarios (e.g., severely low hemoglobin levels and anemia). Our algorithm offers full flexibility, allowing users to choose any copula family and specify different parameters for each bivariate copula in the pair-copula construction.
We simulated scenarios with varying causal coefficients $\beta=\{-0.5,-0.3,0\}$, representing strong, moderate and no effect, respectively and samples sizes $n=\{500,1000,2000\}$. For each combination, we generated 400 datasets. The model $Y=\operatorname{expit}(\alpha+\gamma\,B+\theta\,C-\beta\,\operatorname{cum}(\bar{A}_{4}))$ is fitted using stabilized inverse probability of treatment weights, and an unweighted approach. For each fit, we estimated 95% confidence intervals for $\beta$ using both the robust sandwich estimator and non-parametric bootstrap.
Results are summarized in Table 1, 2 and 3. The inverse-probability-weighted estimator shows minimal bias, with magnitude well below Monte Carlo standard errors. In contrast, the unweighted estimator remains significantly biased across sample sizes, showing no convergence. Both the sandwich and bootstrap CIs of the inverse-probability-weighted estimator approach 95% coverage, though with minor under-coverage. The unweighted estimator, however, exhibits poor coverage due to bias. Statistical power for testing the treatment effect (null hypothesis: $\beta=0$) increases with effect size and sample size for the weighted estimator. In summary, these results confirm that data generated using our simulation algorithm accurately reflect the intended causal effects and time-varying confounding is introduced.
$\beta$ $n$ IPTW Unweighted
--0.5 500 0.027 (0.186) 0.270 (0.115)
1000 0.012 (0.140) 0.269 (0.079)
2000 0.009 (0.096) 0.274 (0.055)
--0.3 500 0.015 (0.179) 0.236 (0.103)
1000 0.004 (0.119) 0.234 (0.073)
2000 0.009 (0.088) 0.235 (0.051)
0 500 0.017 (0.138) 0.189 (0.083)
1000 0.015 (0.087) 0.188 (0.059)
2000 0.014 (0.072) 0.193 (0.043)
Table 1: Bias and Monte Carlo standard errors (in parentheses, defined as $sd(\hat{\beta})$) for inverse-probability-of-treatment-weighted (‘IPTW’) and unweighted estimators (‘unweighted’) across different $\beta$ and $n$ scenarios.
$\beta$ $n$ IPTW Unweighted
Sandwich Bootstrap
$-0.5$ 500 91.8% 93.0% 31.8%
1000 92.0% 91.8% 7.5%
2000 92.3% 92.8% 0.0%
$-0.3$ 500 91.0% 91.5% 35.3%
1000 94.3% 92.5% 8.0%
2000 94.3% 93.8% 0.3%
0.0 500 92.8% 93.0% 41.8%
1000 95.5% 94.3% 12.0%
2000 92.8% 90.0% 0.5%
Table 2: Coverage of 95% confidence intervals for IPTW-sandwich, IPTW-bootstrap, and unweighted estimators across scenarios with varying $\beta$ and $n$ values.
$\beta$ $n$ IPTW Unweighted
Sandwich Bootstrap
$-0.5$ 500 83.3% 79.5% 55.3%
1000 97.5% 97.8% 82.5%
2000 100.0% 100.0% 98.3%
$-0.3$ 500 45.8% 43.0% 12.0%
1000 80.0% 77.8% 15.3%
2000 95.8% 96.5% 26.8%
0.0 500 7.3% 7.0% 58.3%
1000 4.5% 5.8% 88.0%
2000 7.3% 10.0% 99.5%
Table 3: Statistical power to reject the null hypothesis $\beta=0$; that is, no causal effect.
5 Simulating Survival Outcomes
[…]
Figure 2: Example directed acyclic graph showing temporal relationships in a longitudinal survival setting. The sequence of outcomes observed over time is denoted as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\ldots,K+1\}$.
We have described and demonstrated our method to simulate longitudinal data where the outcome is measured at the final time point. Since time-to-event and survival outcomes are very common in medical and epidemiological contexts, we shall now demonstrate how our method can be extended to simulate longitudinal data from survival models.
In Figure 2, each plate represents a single interval. For the $k$th interval, where $k=0,\ldots,K$, the observations include $L_{k},A_{k}$ at the start of the interval, and the survival status $Y_{k+1}$ at the end. The length of each interval is set to 1. In addition to the notation in Section 4, let $T$ denote an individual’s failure time. The outcome is represented be a sequence $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\ldots,K+1\}$ and $Y_{k+1}=I(T<k+1)$ indicates failure.
We describe the algorithm for the case where the time-varying confounder is univariate and no latent processes are present. The algorithm can be extended analogously to the case of ordinary longitudinal outcomes. The illustrative example at the end of this section demonstrates the simulation with two time-varying confounders.
Depending on the chosen marginal structural model, there are various ways to model the outcomes. Examples include:
  • Discrete-time marginal structural model, where the binary survival outcome $Y_{k}$ is modeled directly:
    $P(Y_{k+1}=1\,|\,do(\bar{A}_{k})=\bar{a}_{k},Z=z,Y_{k}=0)=f(\bar{a}_{k},z;\alpha),$ (9)
    where $f$ can be any real-valued functions bounded between 0 and 1, for example, an expit function for a logistic marginal structural model.
  • Continuous-time hazard MSM, where we introduce an auxiliary continuous variable, $\tilde{Y}_{k+1}=T-k\,|\,T>k$, representing the incremental failure time given survival at the start of the $k$th interval. We model incremental survival time $\tilde{Y}_{k+1}$ and truncated it to obtain the observed binary survival outcome $Y_{k+1}$. Given survival at $k$, the time to failure is $T\,|\,T>k=k+\tilde{Y}_{k+1}$. Examples of marginal models for $\tilde{Y}_{k+1}$ include:
    • Cox proportional hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)\exp\{g(\bar{a}_{k},z;\beta)\}$,
    • Additive hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)+g(\bar{a}_{k},z;\gamma)$.
We focus on describing the simulation algorithm for time-to-event outcomes, with a remark on adapting the approach for discrete survival outcomes, which follows a similar but simpler framework.
5.1 Simulation algorithm
The general approach remains similar to the non-survival case, but with the key difference that at each time step $k$, simulation is now conditioned on survival up to the start of the current time interval. For brevity, we continue to omit conditioning on $Z$.
Unlike the derivation of outcome simulation for the non-survival case in Section 4.2, all quantities involved in the copulas must now be conditioned on survival up to this point. Specifically, when simulating the incremental failure time, $\tilde{y}_{k+1}$, given $Y_{k}=0$, (7) becomes:
$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
$=\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j}\}.$ (10)
It is important to note the distinction between the subscripts $j$ and $k$ in the product. To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, which corresponds to the $k$th panel in Figure 2. Each copula in (10) associates $\tilde{y}_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,\ldots,k$.
In the non-survival case discussed in Section 4.2, the quantiles $F(l_{j}\,|\,\bar{l}_{j-1},\bar{a}_{j-1})$, for $j=0,\ldots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation. However, this convenience does not extend when the quantiles are conditioned on survival at later time points. For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim\operatorname{N}(0,1)$. Conditioning on survival after 12 months, however, alters the distribution. The baseline health of the survivors will almost certainly not be representative of the initial population, meaning that the distribution is no longer standard normal. This means that the conditional quantiles of $L_{j}$ that we need are not directly available and require additional computation. Fortunately, employing copulas to define the dependence structure offers an efficient method to derive these conditional distributions, eliminating the need for resource-intensive techniques such as Monte Carlo approximation or numerical integration.
The proposed simulation algorithm follows an iterative structure and is formally presented and validated using an inductive approach.
Proposed simulation method: At time step $k$, we specify the cumulative distribution functions:
$F_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}$ $F_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}$ $F_{\tilde{Y}_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}.$
Additionally, we define $k+1$ bivariate copulas $C_{\tilde{Y}_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$. These copulas captures the dependence between $\tilde{Y}_{k+1}$ and each preceding time-varying covariate. Draw $w_{k,1}$, $w_{k,2}$, $w_{k,3}\sim\operatorname{U}(0,1)$ and compute:
$l_{k}=F^{-1}_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}(w_{k,1}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$ $a_{k}=F^{-1}_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}(w_{k,2}\,|\,\bar{l}_{k},\bar{a}_{k-1}),$ (11)
to obtain a sample $(l_{k},a_{k})$. We initialize by setting $F(\tilde{y}_{k+1}\,|\,\bar{l}_{k},do(\bar{a}_{k}),y_{k}=0)=w_{k,3}$. Then, for $j=k-1,\ldots,0$, we iteratively apply the transformations:
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
$\qquad\quad F(\tilde{y}_{k}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0);\Theta_{{k-1},j}\big{\}}$ (12)
$F(\tilde{y}_{k+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j+1},do(\bar{a}_{k}),y_{k}=0),$
$\qquad\qquad F(l_{j+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j+1}\big{\}}.$ (13)
By induction, the quantiles in the $h$-function in (12) are available from the previous time step $k-1$. We prove this claim as Proposition 5.1.
When the recursion reaches $j=0$, we obtain the quantile $F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0)$. Applying the $h$-function one final time, we set:
$u_{k+1}=F(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}))$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0),$
$\qquad\qquad F(l_{0}\,|\,do(\bar{a}_{k}),y_{k}=0);\Theta_{k,0}\big{\}}.$
A sample of $\tilde{y}_{k+1}$ is then obtained via inversion:
$\tilde{y}_{k+1}=F_{Y_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}^{-1}(u_{k+1}),$ (14)
with the corresponding binary outcome $y_{k+1}=I(\tilde{y}_{k+1}<1)$. This completes the $k$th time step, yielding the sample $(l_{k},a_{k},\tilde{y}_{k+1},y_{k+1})$. If $y_{k+1}=1$, the simulation stops; otherwise, proceed to $k+1$.
Crucially, our method avoids computing the full conditional distribution of $\tilde{Y}_{k+1}\mid\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0$, but rather requiring just a quantile corresponding to our simulated value; this is central to the efficiency of our approach. Equation (12) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$. In (13), we recursively apply $k+1$ inverse $h$-functions to obtain the marginal quantile of $F(\tilde{Y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0)$. Finally, we simulate $\tilde{Y}_{k+1}$ from the target marginal structural model via inversion. In total, this time step requires $42k+$ 4 functions, including the simulation of $L_{k}$ and $A_{k}$.
We proceed to show that the proposed method works using mathematical induction.
Proposition 5.1.
Given the quantiles $w_{k,1},w_{k,2},w_{k,3}$ for each $k=0,\ldots,K-1$ and associated univariate distributions, together with copula parameters $\Theta_{k,j}$ for each $k=0,\ldots,K$ and $j=0,\ldots,k$, we can use the procedure outlined above to simulate a survival outcome according to the specified distributions.
The proof is given in Appendix B.
Remark 5.2.
To simulate a binary survival outcome directly from a discrete marginal structural model, such as that defined in (9), we adopt an approach analogous to Fan et al. (2017). Specifically, we assume that there exists a continuous latent random variable, $S_{k+1}$, whose distribution under the intervention $do(\bar{A}_{k})$ is $F_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}$, such that the binary survival indicator is obtained by dichotomizing $S_{k+1}$ at a constant threshold $C$. This formulation allows dependencies between $Y_{k+1}$ and past covariates to be modelled via the latent variable $S_{k+1}$, and the simulation procedure follows the algorithm described above, replacing $\tilde{Y}_{k+1}$ by $S_{k+1}$. Given the marginal quantile $u_{k+1}$, inversion yields $s_{k+1}=F^{-1}_{S_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),y_{k}=0)$ Finally, the binary outcome is obtained as $y_{k+1}=I(s_{k+1}>C)$.
5.2 Extensions and computational complexity
In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(2k+4)=K^{2}+3K$ calculations. This implies that the computational complexity grows quadratically with $K$. However, if an individual experiences failure at any time step, the simulation terminates early, meaning that this complexity represents the maximum possible computational cost per individual.
A key extension involves simulating multiple outcomes, and the approach depends on whether the additional outcomes pertain to survival or non-survival events. For non-survival outcomes simulated from marginal structural models—as opposed to conditional distributions—the number of calculations to derive $F(\bm{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y^{\triangle}_{k}=0)$ increases proportionally with the number of outcomes, $m$. Here, $\bm{Y}_{k+1}$ is an $m$-dimensional vector, and $Y_{k}^{\triangle}$ denotes the primary survival outcome, with all other outcomes assumed to be non-survival. To incorporate dependence between outcomes given past variables, we must specify the joint distribution $\bm{Y}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k}$ and simulate from this using a pair-copula construction. The total number of calculations is $\sum_{k=0}^{K-1}\big{\{}(k+3)m+k+1\big{\}}=(K+5)Km/2+K(K+1)/2$, which scales linearly with $m$.
For competing risks, such as death from other causes, additional calculations are needed to determine the distribution of $L$ given survival from all risks, as well as its distribution conditioned on the occurrence of each specific risk. Consequently, the computational cost scales proportionally with $m$. Furthermore, iterative evaluations of inverse $h$-functions are required to obtain marginal interventional quantiles for each risk. The total number of calculations is given by $\sum_{k=0}^{K-1}(2mk+3m+1)=(K^{2}+2K)m+K$ which grows linearly with $m$. The simulation algorithm for a setting with a single competing risk is described in Appendix C.
Another important extension involves including multiple time-varying confounders, observed or unobserved, such that $\bm{L}$ becomes a vector of $p$ variables. The number of calculations required for the distributions of ¯$\bm{\bar{}}L_{k-1}$, and $F(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)$ among survivors at $k$ scales proportionally with $p$. The total number of computations becomes $\sum_{k=0}^{K-1}(2pk+3p+2)=K(K+2)p+2K$. Thus, the computational complexity grows linearly with $p$. In the example presented in Section 5.3, we simulate data with two time-varying confounders.
All of these calculations require evaluations of $h$-functions and their inverses. For commonly used copulas with analytic $h$-functions, these evaluations are computationally efficient and fast.
5.3 Illustrative example
We use an example to demonstrate the proposed algorithm simulating longitudinal time-to-event data, following a setup similar to Seaman and Keogh (2024). We consider baseline variables $B$ and $C$, two time-varying covariates $L_{k}$ and $W_{k}$, and treatment $A_{k}$ at follow-up $k$. The time-to-event is $T$, with $Y_{k+1}=I(T<k+1)$ indicating failure status observed at the end of the $k$th interval. The auxiliary continuous variable $\tilde{Y}_{k+1}$ is defined as previously. The data-generating models for the baseline and time-varying covariates, and treatment status are detailed in Appendix D. We assume a Cox marginal structural model given by:
$\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),c,y_{k}=0)=\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right),$ (15)
where $\lambda_{0}=\exp(-2)$ is the baseline hazard. We set $\beta_{C}=-0.5$ and vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$. This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit. This implies that given survival at follow-up time $k$, the incremental survival time $\tilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
$\tilde{Y}_{k+1}\sim\operatorname{Exp}(\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right)).$
If $\tilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise. We set the total number of follow-ups $K$ to 6. We set the copulas describing the association between $L_{j}$ ($j=0,\ldots,k$) and $Y_{k+1}$ to be Student t with parameter $\rho_{L}$ and 5 degrees of freedom, although each pairwise dependence could, in principle, be distinct. Similarly, we model the copulas capturing the dependence between $Y_{k+1}$ and each $W_{j}$ ($j=0,\ldots,k$), given $L_{k}$, as Gaussian with correlation $\rho_{W}$, though again we note that different copulas could be specified for each time point. We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding (see Table 6 in Appendix D). We simulate data for sample size $n=\{1000,2000,5000\}$.
Remark 5.3.
Table 7 (in Appendix D) reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario. The simulations were carried out on a laptop using a single score without parallelization. Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute. This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values.
We contrast two approaches for fitting the Cox marginal structural model specified in (15): stabilized inverse-probability-weighting and unweighted estimation. Details on weight calculation are provided in Appendix D. Table 4 summarizes the bias of estimates from 500 replications, compared to corresponding Monte Carlo standard errors as a measure of uncertainty. As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard errors. In contrast, the unweighted estimator exhibits considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, which persists even at $n=5000$. The magnitude of this bias far exceeds the Monte Carlo standard errors. The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $-0.04(2.27)$ $-1.05(1.54)$ $-0.14(2.31)$ $3.73(1.73)$ $0.27(3.09)$ $0.49(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.68)$ $-1.03(1.06)$ $-0.02(1.63)$ $3.81(1.21)$ $0.02(2.09)$ $0.34(1.51)$ $5000$ $0.02(1.25)$ $-0.96(0.68)$ $0.10(1.08)$ $3.89(0.76)$ $-0.09(1.58)$ $0.24(0.97)$ $1000$ $-0.03(2.35)$ $-0.93(1.59)$ $0.02(1.99)$ $3.64(1.46)$ $0.03(2.78)$ $0.44(1.94)$ $\beta_{A}=0.2$ $2000$ $-0.11(1.60)$ $-0.93(1.11)$ $0.02(1.49)$ $3.65(1.18)$ $0.17(2.02)$ $0.49(1.51)$ $5000$ $-0.03(1.14)$ $-0.95(0.71)$ $0.04(1.01)$ $3.65(0.71)$ $0.07(1.36)$ $0.51(0.89)$ Medium $1000$ $-0.07(2.04)$ $-0.43(1.51)$ $-0.13(2.19)$ $1.90(1.73)$ $0.23(2.91)$ $0.08(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.02(1.60)$ $-0.35(1.01)$ $0.00(1.65)$ $1.95(1.22)$ $0.09(2.27)$ $-0.01(1.55)$ $5000$ $-0.09(1.10)$ $-0.42(0.64)$ $-0.01(1.21)$ $1.97(0.80)$ $0.07(1.59)$ $-0.02(1.03)$ $1000$ $-0.12(2.06)$ $-0.36(1.43)$ $-0.04(2.02)$ $1.75(1.54)$ $0.14(2.70)$ $0.09(1.92)$ $\beta_{A}=0.2$ $2000$ $-0.10(1.62)$ $-0.42(1.08)$ $0.03(1.49)$ $1.68(1.11)$ $0.09(2.09)$ $0.17(1.44)$ $5000$ $-0.03(1.04)$ $-0.39(0.66)$ $0.03(0.95)$ $1.78(0.68)$ $0.03(1.32)$ $0.14(0.88)$ Low $1000$ $0.01(1.82)$ $0.46(1.29)$ $-0.23(2.36)$ $-1.95(1.83)$ $0.21(2.90)$ $-0.20(2.20)$ $\beta_{A}=-0.5$ $2000$ $-0.10(1.32)$ $0.44(0.89)$ $-0.04(1.69)$ $-1.83(1.25)$ $0.07(2.13)$ $-0.30(1.59)$ $5000$ $-0.02(1.04)$ $0.47(0.58)$ $0.01(1.07)$ $-1.77(0.84)$ $0.03(1.46)$ $-0.32(0.99)$ $1000$ $0.04(1.98)$ $0.53(1.30)$ $0.05(1.98)$ $-1.88(1.46)$ $-0.03(2.65)$ $-0.31(1.88)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.45)$ $0.43(0.99)$ $0.04(1.43)$ $-1.85(1.10)$ $0.01(1.87)$ $-0.26(1.38)$ $5000$ $-0.01(0.92)$ $0.43(0.59)$ $0.07(0.95)$ $-1.89(0.68)$ $-0.06(1.25)$ $-0.25(0.84)$
Table 4: Bias and Monte Carlo standard errors (×$\times 10$) of estimates produced by inverse probability of treatment weighted (‘IPTW’) and unweighted Cox regressions, under different strengths of confounding, effect sizes, and sample sizes.
Table 5 presents coverage rates of the 95% confidence intervals. Confidence intervals for the weighted estimated, computed using both sandwich variance estimation and bootstrap resampling, achieve coverage close to the nominal level, with slight undercoverage. In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, especially for the treatment effect estimate, $\beta_{A}$.
Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.8% 95.2% 88.4% 93.4% 92.8% 42.2% 93.8% 93.8% 93.4% $\beta_{A}=-0.5$ 2000 94.0% 94.0% 84.4% 95.2% 95.6% 12.4% 95.6% 95.6% 94.6% 5000 93.2% 92.8% 70.2% 94.4% 93.4% 0.2% 94.2% 93.6% 93.8% 1000 92.8% 93.8% 91.8% 96.2% 95.8% 31.6% 93.4% 94.0% 95.6% $\beta_{A}=0.2$ 2000 95.2% 95.0% 87.6% 94.6% 95.0% 9.8% 94.8% 94.6% 92.2% 5000 93.6% 93.6% 72.4% 93.8% 93.0% 0.2% 94.0% 94.0% 90.0% Medium 1000 93.8% 94.2% 93.4% 94.8% 93.6% 79.6% 93.8% 94.0% 93.6% $\beta_{A}=-0.5$ 2000 93.0% 93.8% 92.4% 95.2% 94.6% 64.8% 92.2% 93.4% 96.2% 5000 95.4% 94.8% 90.0% 93.2% 92.6% 30.2% 93.0% 93.8% 93.6% 1000 94.2% 94.8% 95.0% 95.0% 93.8% 77.2% 94.2% 93.2% 95.8% $\beta_{A}=0.2$ 2000 94.6% 93.8% 91.6% 93.6% 93.2% 64.0% 92.8% 92.2% 94.2% 5000 95.4% 94.6% 91.6% 95.4% 95.2% 26.8% 94.8% 94.4% 95.4% Low 1000 95.4% 95.4% 93.0% 95.0% 95.2% 84.0% 95.2% 95.2% 94.6% $\beta_{A}=-0.5$ 2000 95.6% 96.4% 93.0% 95.8% 96.0% 71.6% 95.0% 95.6% 94.6% 5000 93.2% 93.2% 87.2% 94.2% 94.8% 42.6% 94.0% 94.4% 94.4% 1000 94.6% 95.6% 94.8% 96.4% 96.6% 77.4% 92.8% 94.8% 96.0% $\beta_{A}=0.2$ 2000 94.8% 95.6% 92.8% 94.4% 94.6% 60.6% 95.0% 95.2% 94.2% 5000 94.8% 94.2% 89.4% 95.0% 95.4% 20.0% 95.0% 94.2% 93.8%
Table 5: Coverage rates for IPTW-Sandwich (W-S), IPTW-Bootstrap (W-B), and unweighted (UW) estimators of $\beta_{C}$, $\beta_{A}$, and $\beta_{AC}$ under different strengths of confounding, effect sizes, and sample sizes.
In summary, the asymptotically unbiased nature of the inverse probability weighted estimates, compared to the biased unweighted estimates, further confirms that the simulated data were generated in accordance with the Cox model in (15), and that confounding by $L$ and $W$ was appropriately introduced.
6 Discussion
In this paper, we proposed and demonstrated a principled strategy for simulating longitudinal data exactly from marginal structural models, leveraging the frugal parameterization (Evans and Didelez, 2024). Our proposed method is highly flexible and can accommodate a variety of practical scenarios, including multiple treatments (for example, primary and drop-in), dynamic treatment regimes, dose-response analyses, multiple outcomes (such as secondary outcomes and competing risks), and censoring. Additionally, it imposes no restrictions on the data-generating process or marginal structural causal models. In Section 5, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials. Our methods are both conceptually and computationally simple, which crucial for practical applications such as sample size calculation.
[…]
Appendix A Extension to Include Time-Varying Latent Variables
We provide an example of how to incorporate a latent process that confounds the covariates and the final outcome following the algorithm outlined in Section 4. The steps to simulate longitudinal data from the model depicted in Figure 3 are as follows:
[…]
Figure 3: Extension of the model in Figure 1 which allows for an unobserved process $H_{k}$ affecting only $Z$, $L_{k}$ and $Y$.
  1. 1.
    Baseline simulation: The directed acyclic graph in Figure 3 indicates that the latent variable $H_{0}$ precedes $Z$ in topological order, potentially indicating an initial selection bias. Accordingly, we simulate baseline values $(z,h_{0},l_{0},a_{0})$ sequentially:
    $h_{0}$ $=F_{H_{0}}^{-1}(w_{0,0})$
    $z$ $=F_{Z\mid H_{0}}^{-1}(w_{Z}\,|\,h_{0})$
    $l_{0}$ $=F_{L_{0}\mid ZH_{0}}^{-1}(w_{0,1}\,|\,z,h_{0})$
    $a_{0}$ $=F_{A_{0}\mid ZH_{0}L_{0}}^{-1}(w_{0,2}\,|\,z,h_{0},l_{0}),$
    where $w_{Z},w_{0,0},w_{0,1},w_{0,2}$ are standard uniform variables.
  2. 2.
    Time-step simulation: For each time step $k=1,\dots,K$, simulate:
    $h_{k}$ $=F_{H_{k}\mid\bar{H}_{k-1}}^{-1}(w_{k,1}\,|\,\bar{h}_{k-1})$
    $l_{k}$ $=F_{L_{k}\mid Z\bar{H}_{k}\bar{L}_{k-1}\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{h}_{k},\bar{l}_{k-1},\bar{a}_{k-1})$
    $a_{k}$ $=F_{A_{k}\mid Z\bar{H}_{k}\bar{L}_{k}\bar{A}_{k-1}}^{-1}(w_{k,3},\,|\,z,\bar{h}_{k},\bar{l}_{k},\bar{a}_{k-1})$
    where $w_{k,1},w_{k,2},w_{k,3}\sim U(0,1)$.
  3. 3.
    Outcome simulation: Once all time-varying covariates and treatment have been simulated, we apply the same derivation following the frugal parameterization as in Section 4.2 and simulate $Y$ in two steps:
    1. (a)
      Compute the marginal quantile:
      $u^{*}_{Y}$ $=h^{-1}\big{[}\dots h^{-1}[h^{-1}\left\{h^{-1}\left(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{h}_{K},z,\bar{a}_{K-1});\Theta_{K,1}\right),F(h_{K}\,|\,\bar{l}_{K-1},\bar{h}_{K-1},z,\bar{a}_{K-1});\Theta_{K,2}\right\}$
      $F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{h}_{K-1},z,\bar{a}_{K-2});\Theta_{K-1,1}]\dots;\Theta_{0,2}\big{]}$
      where $w_{Y}\sim\operatorname{U}(0,1)$;
    2. (b)
      Obtain the outcome: $y=F_{Y\mid Z,\bar{A}_{K}}^{*{-1}}(u^{*}_{Y}\,|\,z,do(\bar{a}_{K}))$.
Compared to the case where without latent variable, the number of applications of inverse $h$-functions in Step 3(a) doubles. These transformations systematically disentangle the dependencies between $Y$ and $H_{k}$, as well as between $Y$ and $L_{k}$. In addition to the information required in the case without latent confounders, the user also specifies:
  • the development of the latent confounder $H_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k}$;
  • $K+1$ pair-copula densities describing dependencies between the outcome and each member of the latent process: $c_{YH_{k}\mid Z,\bar{H}_{k-1},\bar{L}_{k-1},do(\bar{A}_{k-1})}$ for $k=0,1,\ldots,K$.
Extending the framework to include additional time-varying covariates follows a similar approach. In essence, simulating a latent process is technically equivalent to simulating any other time-varying confounder. Whether a confounder is observed or unobserved does not alter the fundamental simulation methodology. With common choices of parametric bivariate copulas, evaluating the $h$-functions is fast, making the simulation efficient. As the dimensionality of time-varying confounders, either observed or unobserved, the computational complexity grows linearly.
Appendix B Proof of Proposition 5.1
Proof.
We prove that the algorithm correctly simulates data for all $k\geq 0$.
Base case ($k=0$):
Initialize the algorithm by specifying the cumulative distribution functions $F_{L_{0}}$, $F_{A_{0}\mid L_{0}}$, $F_{\tilde{Y}_{1}\mid do(A_{0})}$, and the copula $C_{\tilde{Y}_{1}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{0,0}$. Draw $w_{0,1}$, $w_{0,2}$, $w_{0,3}\sim U(0,1)$ and compute:
$l_{0}=F^{-1}_{L_{0}}(w_{0,1})$ $a_{0}=F^{-1}_{A_{0}\mid L_{0}}(w_{0,2}\,|\,l_{0}),$ (16)
to obtain a sample $(l_{0},a_{0})$. Using the transformation:
$F_{\tilde{Y}_{1}\,|\,do(A_{0})}(\tilde{y}_{1}\,|\,do(a_{0}))$ $=h^{-1}\big{\{}F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0})),F_{L_{0}}(l_{0});\theta_{0,0}\big{\}},$ (17)
set $w_{0,3}=F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0}))$ and substitute $F_{L_{0}}(l_{0})$ with $w_{0,1}$, yielding:
$u_{1}=h^{-1}\left(w_{0,3},w_{0,1};\theta_{0,0}\right).$ (18)
By inversion, compute $\tilde{y}_{1}=F_{Y_{1}\mid do(A_{0})}^{-1}(u_{1})$, and the corresponding binary outcome $y_{1}=I(\tilde{y}_{1}<1)$. This completes the first iteration, yielding a sample $(l_{0},a_{0},\tilde{y}_{1},y_{1})$. If $y_{1}=1$, stop; otherwise, proceed to $k=1$.
For $k=1$, define the necessary cumulative distribution functions $F_{L_{1}\,|\,L_{0},A_{0}}$, $F_{A_{1}\mid\bar{L}_{1},A_{0}}$, $F_{\tilde{Y}_{2}\mid do(\bar{A}_{1})}$, and copulas $C_{\tilde{Y}_{2}L_{1}\mid L_{0}do(\bar{A}_{1})}$ and $C_{\tilde{Y}_{2}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{1,1}$ and $\theta_{1,0}$, respectively. These two copulas associate $\tilde{Y}_{2}$ with previous time-varying confounders. Draw $w_{1,1}$, $w_{1,2}$, $w_{1,3}$ from $U(0,1)$, and obtain $(l_{1},a_{1})$ as before.
We now need to obtain a sample of $\tilde{y}_{2}$. Calculate:
$F(l_{0}\,|\,do(a_{0}),y_{1}=0)$ $=h\big{\{}F(l_{0}),F(\tilde{y}_{1}\,|\,do(a_{0}));\theta_{0,0}\big{\}}$ (19)
$F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,\bar{l}_{1},do(\bar{a}_{1}),y_{1}=0),F(l_{1}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0);\theta_{1,1}\big{\}}$
$F(\tilde{y}_{2}\,|\,do(a_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0),F(l_{0}\,|\,do(\bar{a}_{1}),y_{1}=0);\theta_{1,0}\big{\}}.$
Note that we omit some subscripts on quantiles for brevity. In (19) we derive the quantile of $L_{0}$ given $a_{0}$ among survivors the start of the $k=1$ interval, where $F(l_{0})=w_{0,0}$ and $F(\tilde{y}_{1}\,|\,do(a_{0})$ is already computed in (18) in step $k=0$. All other necessary quantities are specified at $k=1$. Therefore, the base case is satisfied since the algorithm simulates the first time step and provides the required inputs for the next.
Inductive step:
Assume the algorithm correctly simulates up to the $k$th iteration. We have $F(l_{k}\mid\overline{l}_{k-1},do(\bar{a}_{k-1}),y_{k}=0)$ and $F(y_{k+1}\mid\overline{l}_{k},do(\bar{a}_{k}),y_{k}=0)$, as well as a collection of pair copulas between $Y_{k+1}$ and $L_{j}$ given $(\bar{l}_{j-1},\bar{a}_{j-1})$ with parameters $\theta_{k,j}$. We need to compute the updated quantiles $F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$ for each $j<k$, given from the previous iteration:
$F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0),$ $F(y_{k}\mid\overline{l}_{j},do(\bar{a}_{k-1}),y_{k-1}=0).$
This can be achieved by noting that
$F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k-1}),y_{k-1}=0)=F(l_{j}\mid\overline{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0),$
which allows us to apply (12) to compute the required quantiles.
We must then obtain $F(y_{k+1}\mid\overline{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ for each $j$ to use in the next iteration. This can be done using (13), and hence, we can move to the $(k+1)$th iteration.
By the principle of induction, the statement is true for all $k\geq 0$. ∎
Appendix C Competing risk
At each time step $k$, after simulating time-varying covariate $L_{k}$ and treatment $A_{k}$, we generate the primary outcome $Y_{k}$ and a competing risk $D_{k}$ using their respective auxiliary variables $\tilde{Y}_{k}$ and $\tilde{D}_{k}$, following the approach outlined in Section 5. For brevity, baseline covariates $\bm{Z}$ continue to be omitted, but in settings where they are present, they can be simulated as the initial step and included in the conditioning set of all subsequent quantities.
Under this simulation framework, we specify:
  • A marginal structural model for the primary event: $F^{*}(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
  • A marginal structural model for the competing event: $F^{*}(\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
  • A bivariate copula parameterized by $\Psi_{k+1}$ to model the dependence between the above interventional margins of $\tilde{Y}_{k+1}$ and $\tilde{D}_{k+1}$;
  • Bivariate copulas that ‘weave’ the multivariate dependence between $Y_{k+1}$ and past covariates $\bar{L}_{k}$, given $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$;
  • Bivariate copulas that ‘weave’ the multivariate dependence between $\tilde{D}_{k+1}$ and $\bar{L}_{k}$ conditional on $\tilde{Y}_{k+1}$, $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Phi_{k,j}$ for $j=0,\ldots,k$.
The joint outcome model $p(\tilde{y}_{k+1},\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$ factorizes as follows:
Primary event model:
$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$
$=p^{*}(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)\times$
$\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0,D_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\}.$ (20)
Competing event model:
$p(\tilde{d}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)$
$=p(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)\times$
$\prod_{j=0}^{k}c^{*}_{\tilde{D}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},\tilde{Y}_{k+1}D_{k}=0}\big{\{}F(d_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0);\Phi{k,j}\}.$ (21)
The first term in (21), the distribution of $\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0$ can be obtained through:
$F(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)=h\big{\{}F*(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0),F^{*}(\tilde{y}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)\big{\}}.$
The challenge lies in computing $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ in (20), and $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ in (21), for $j=0,\ldots,k$. The former can be calculated in the same iterative fashion as described in Section 5. The latter can be derived as:
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
$\qquad F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\big{\}}.$ (22)
Notably, the two components in the $h$-function in (22) exactly correspond to the those in the primary event model in (20), except with the order of conditioning swapped. Since these values are computed during the simulation of $\tilde{y}_{k+1}$, requiring only $k+1$ additional evaluations of the $h$-functions in (22).
The algorithm for simulating each time step $k$ for an individual can be summarized into:
  1. 1.
    Simulate the time-varying confounder $L_{k}$ and treatment $A_{k}$ from their respective distributions.
  2. 2.
    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=0,\dots,k-1$.
  3. 3.
    Compute $F(y_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=1,\dots k$, leading to $u_{k+1}=F(\tilde{y}_{k+1}\,|\,z,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
  4. 4.
    Compute: $\tilde{y}_{k+1}=F^{*-1}_{\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
  5. 5.
    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ for $j=0,\ldots,k$ by evaluating $k+1$ $h$-functions following (22).
  6. 6.
    Compute $F(\tilde{d}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1}=0,D_{k}=0)$ for $j=1,\ldots k$, leading to $v_{k+1}=F(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
  7. 7.
    Compute: $\tilde{d}_{k+1}=F_{\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),\tilde{Y}_{k+1},D_{k}=0}^{*-1}(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
This approach requires at most $74k+$ 7 evaluations of analytic functions per individual per time step. An example implementation is available at https://github.com/XiLinStats/LongSiMSM.
Appendix D Additional details of the illustration in Section 5.3
Models for baseline, time-varying covariates and treatment
A baseline:
$B\sim\operatorname{Exp}(2)$ $L_{0}\sim\operatorname{Gamma}(k=1+0.5\,B+0.5\,C,\;\theta=1)$
$C\sim\operatorname{Bernoulli}(0.7)$ $W_{0}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C\}).$
We specify that throughout the treatment assignment is influenced by $B$, $C$, $L_{k}$ and $W_{k}$ as:
$A_{k}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-1.5+0.5\,B+0.5\,C+0.5\,L_{k}+0.5\,W_{k}\}).$
For subsequent follow-up visits $k>0$:
$L_{k}$ $\sim\operatorname{Gamma}(k=1+0.5\,B+0.5\,C+0.1\,L_{k-1}-0.5\,A_{k-1},\;\theta=1)$
$W_{k}$ $\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C+W_{k-1}-0.6\,A_{k-1}\}).$
Parameter Scenarios
Table6 summarizes the parameter values for treatment effect size and confounding scenarios:
Effect size Large $(\beta_{A},\beta_{CA})=(-0.5,0.3)$ Small $(\beta_{A},\beta_{CA})=(0.2,0.1)$ Level of confounding High $(\rho_{L},\rho_{W})=(0.9,0.7)$ Medium $(\rho_{L},\rho_{W})=(-0.5,0.4)$ Low $(\rho_{L},\rho_{W})=(-0.1,-0.2)$
Table 6: Parameter values of treatment effect size and confounding scenarios.
Runtime details
Table 7 reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario. The simulations were carried out on a local laptop using a single score without parallelization. Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute. This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values. For example, determining the sample size might require testing 10 different parameter combinations (such as effect sizes, prevalence of drop-in treatments, or adherence levels) across four sample sizes, each bootstrapped 200 times to estimate statistical power. This requires $10\times 4\times 200=8,\!000$ simulations. Using our method, suppose $n=10,000$ and $K=10$, the total runtime is approximately 13 hours on a single core, which can be reduced to about 2 hours using an 8-core laptop or under 10 minutes on a 50-core CPU through parallelization.
In contrast, simulation methods requiring computationally intensive operations, such as Monte Carlo approximations (see, for example, Seaman and Keogh, 2024) or numerical integration, are often too slow to support such large-scale simulations. This highlights a key advantage of our proposed method: its ability to handle extensive simulation demands efficiently, making it well-suited for practical applications.
$n$ $K=5$ $K=10$ 1,000 0.27 0.79 2,000 0.45 1.68 5,000 1.69 2.49 10,000 2.70 6.09 50,000 16.50 45.49
Table 7: Runtime (in seconds) to simulate a dataset of size $n$, averaged over 20 repeats.
Weight calculation
To fit the Cox marginal model, we employ stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
$\prod_{j=0}^{k}\frac{\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)}{\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)},$
where $\hat{p}_{j}(A_{j}\mid C,\bar{A}_{j-1},Y_{j}=0)$ and $\hat{p}_{j}(A_{j}\mid B,C,\bar{L}_{j},\bar{W}_{j},\bar{A}_{j-1},Y_{j}=0)$ are estimated using logistic regression models fitted to the survivors at time step $j$.
The denominator in the weight formula represents the correctly specified propensity score model under the assumed data-generating process, while the numerator stabilizes the weights. To avoid unintended confounding, the baseline covariates included in the numerator must also be adjusted for in the marginal structural model (Cole and Hernán, 2008). In accordance with the model in (15), we only include $C$ in the numerator for stabilization.